green09:

Timmy and Jerod are eating cupcakes. If they each can eat 2 cupcakes to 3 minutes, how long will it take the two of them to eat 18 cupcakes

justjm:

It's not 20 There are various ways to solve this and one of the ways is to form a proportion. You can do cupcakes / minute \[\frac{2~\text{cupcakes}}{3~\text{minutes}} = \frac{18~\text{cupcakes}}{x~\text{minutes}} \] Now you can cross multiply to get \(2x=18(3)\) and isolate x to get the time it takes for the two of them to eat 18 cupcakes.